INFRARED SPECTROSCOPY OF Li(METHYLAMINE)n(NH_{3})m CLUSTERS

\underline{N. BHALLA}, L. VARRIALE, N. M. TONGE AND A. M. ELLIS, Department of Chemistry, University of Leicester, Leicester, LE1 7RH, United Kingdom..

Recent work in our laboratory has led to the first size-selective infrared spectra of neutral Li(NH_{3})n clusters. Here we report on new work extending these studies to mixed cluster species, where ammonia derivatives such as methylamine (MA) replace some of the ammonia molecules. Spectra recorded using mass-selective IR depletion spectroscopy will be reported for Li(MA)m(NH_{3})n, with the emphasis on n + m = 4 clusters. The impact of replacing NH_{3} with bulkier ligands on the solvent structure will be described. In addition the IR spectrum of the smaller cluster Li(MA)(NH_{3}) will be reported for the first time.